# The thermodynamic uncertainty relation of a quantum-mechanically coupled two-qubit system

### v1

@@ Abstract
The minimal bound of the thermodynamic uncertainty relation (TUR) is modulated from that of the classical counterpart ($\mathcal{Q}_{\rm min}=2$) when a quantumness is present in the dynamical process far from equilibrium. A recent study on a dissipative two-level system (TLS) subject to an external field indicates that quantum coherences can suppress the fluctuations of the irreversible current and loosens the TUR bound to $\mathcal{Q}_{\rm min}^{\rm TLS}\approx 1.25$. Here, we extend on the field-driven single TLS to a quantum-mechanically coupled two-qubit system (TQS), and explore, in addition to the single qubit coherence, how the quantum coupling between the two qubits affects the photon current, fluctuations, and the TUR bound. We find that the TUR bound of TQS depends on the strength of coupling, such that $\mathcal{Q}_{\rm min}^{\rm TQS}=\mathcal{Q}_{\rm min}^{\rm TLS}\approx 1.25$ when the two qubits are effectively decoupled under weak coupling, whereas another loose bound $\mathcal{Q}_{\rm min}^{\rm TQS}\approx 1.36$ is identified for two strongly coupled qubits under strong fields. By contrasting the TQS against two coupled noisy oscillators, we illuminate the quantumness unique to the TQS and its effect on the TUR.
@@
Since its first proposal [1] and subsequent proof for continuous time Markov jump processes on networks and overdamped Langevin dynamics [2, 3], the thermodynamic uncertainty relation (TUR) has been extended to more general contexts along with its variations over the past decade [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The TURs are formulated in the form of a trade-off relation between dissipation and precision of a dynamical observable along with its minimal bound. The relations can be considered as a quantitative version of thermodynamic principle for nonequilibrium processes that constrains the total entropy production more tightly than the second law as $\Delta S_{tot}(t)/k_{B}\geq 2\langle\theta(t)\rangle^{2}/\langle\delta\theta^{2}(t)\rangle$, where $\theta(t)$ denotes a current-like observable with odd parity upon time reversal $\theta(-t)=-\theta(t)$ [7, 23].
@@
Several studies that examined the TURs for open quantum processes have indicated a loosening of its minimal bound, leading to the TUR violation [24, 25, 26, 27, 28, 29, 30, 31], signifying that quantumness can enhance the precision of the dynamical processes beyond its classical limit set by the conventional TUR. Recent works on a field-driven two-level system (TLS) and three-level $\Lambda$-system immersed in photonic bath [32, 33] have shown that the size of the fluctuations in photon current is decided by the competition between the real and imaginary parts of the quantum coherence that arises from a superposition between different states in a given basis. The imaginary part of coherence, which was shown to be responsible for the generation of non-equilibrium current [34, 35], is linked to the dissipative flow of energy or absorption between states that reduces the current fluctuations. In contrast, the real part, related to the dephasing mechanism accompanied by population mixing, amplifies the fluctuations [32, 33].
@@
Here, we consider a theoretical model of quantum-mechanically coupled two-qubit system (TQS), in which two atoms with ground and excited states [36] are in a photonic bath while an external field continuously irradiates only one of the two atoms. The dynamics in the first atom can affect the quantum state of the second atom due to the coupling, or it simply emits and dissipates a photon into the bath, the process of which can be formulated with the corresponding Lindblad equations.
@@
We characterize non-equilibrium responses of the TQS by means of the uncertainty product of TUR ($\mathcal{Q}$), finding that the maximal precision in the photon current is attained when both the strengths of field and two-qubit coupling are moderate. Our study offers physical insights into this finding by discussing the details of how the coherence, correlation, and entanglement of TQS arise from the varying strength of external field and quantum coupling of the two qubits, and contributes to its TUR.
@@ II.1 Two-qubit system subject to an external field and its evolution equation
where $K$ is the coupling strength. The external field with the frequency $\omega$ irradiates the 1st qubit, resulting in an interaction energy Hamiltonian
@@
where the dipole operator $\vec{d}=\vec{d}_{10}\sigma^{1}_{+}+\vec{d}_{01}\sigma^{1}_{-}$ acts on the first qubit and the electric field is given by $\vec{E}(r,t)\simeq\vec{\epsilon}e^{i\omega t}+\vec{\epsilon}^{\ast}e^{-i\omega t}$. Thus, the 1st qubit display unitary oscillation between the ground and the excited states with the Rabi frequency $\Omega=\vec{d}_{10}\cdot\vec{\epsilon}/\hbar=\vec{d}_{01}\cdot\vec{\epsilon}^{\ast}/\hbar$. The Hamiltonian for the surrounding bath is
@@
The Lindblad equation that casts the system-bath interaction in a specific form is obtained by tracing out the bath degrees of freedom and approximating the hierarchical equation under the weak-system bath coupling [37]. The reduced density operator for the system $\rho(t)$ undergoes a dynamic evolution, obeying the following equation along with the Lindblad dissipator, $\mathcal{D}\left(\rho(t)\right)$, working on both the qubits
@@
for $i=1,2$. Here, the term $\bar{n}_{i}=(e^{\beta\hbar\omega_{i}}-1)^{-1}$ denotes the mean occupation number of the bath affected by the dynamics of the $i$-th qubit, $\{A,B\}=AB+BA$ is the anti-commutator, and $\gamma_{1,2}$ is the relaxation rate of the qubit to its ground state caused by the interaction with the bath. The condition of the weak system-bath coupling is dictated by the condition of $\omega\gg\gamma_{1},\gamma_{2}$. Rescaling the time scale by $\gamma_{1}$ as $t\rightarrow\tau=\gamma_{1}t$, and all the frequencies discussed here as $\gamma_{2}\rightarrow\gamma\equiv\gamma_{2}/\gamma_{1}$, $\Omega\rightarrow\Omega/\gamma_{1}$ yield a set of coupled equations for the density matrix elements.
@@
After applying the rotating-wave approximation (RWA) with $\omega\gg\Omega$ and representing each density matrix element in a rotating frame, we obtain the Liouville equation for the vectorized form of the reduced density matrix in the Fock-Liouville space,
@@ II.3 Uncertainty product for TQS
Using the definition of photon current through the $i$-th qubit $\langle j_{i}\rangle=\langle n_{i}(\tau)\rangle/\tau$, one can recast Eq. (14) into the product between the affinity ($\mathcal{A}=\Delta S_{tot}/(k_{B}\langle n(\tau)\rangle)$) and the Fano factor ($\mathcal{F}$) of photon current:
@@
where $j=\sum_{i=1,2}j_{i}$. Here, we use the fact that the total entropy production from the system is contributed by the net number of photon emissions that have occurred in both qubits, namely, $\Sigma^{\rm ss}\tau=\mathcal{A}_{1}\langle n_{1}(\tau)\rangle+\mathcal{A}_{2}\langle n_{2}(\tau)\rangle$, with $\mathcal{A}_{i}=\hbar\omega_{i}/k_{B}T$, which thus yields $\mathcal{A}=\frac{\sum_{i=1,2}\mathcal{A}_{i}\langle j_{i}\rangle}{\sum_{i=1,2}\langle j_{i}\rangle}$. We note that in order to be consistent with the RWA, the condition of $\omega\gg\delta\omega_{i}$ is essential. Consequently, $\mathcal{A}_{i}=\hbar(\omega+\delta\omega_{i})/k_{B}T\approx\hbar\omega/k_{B}T$, and we assume hereafter that the two qubits have the same affinity, $\mathcal{A}_{1}\approx\mathcal{A}_{2}$. The mean current $(\left\langle j\right\rangle)$ and the current fluctuations $({\rm Var}(j))$ at steady states are calculated using the method of cumulant generating function (see Methods).
@@ III.1 Populations and coherences of TQS and current-coherence relations
Fig. 2 demonstrates the coherences and currents of the TQS under varying amount of detunings ($\delta\omega_{1}$, $\delta\omega_{2}$) for select values of Rabi frequency ($\Omega/\Omega_{o}=0.1$, 1, 10) and coupling strength ($K=0.1,1,10$). For the sake of our discussion, we normalize the Rabi frequency by $\Omega_{o}=0.37$ that minimizes $\mathcal{Q}^{\rm TLS}$ (Eq. (22)) [32], defining $\tilde{\Omega}\equiv\Omega/\Omega_{o}$. As commented in Eq. (21), the photon current generated in an isolated qubit interacting with an external field is directly proportional to the imaginary part of coherence. This situation arises when the 1st qubit is subject to the external field, whereas the 2nd qubit is effectively decoupled from the 1st qubit due to the weak coupling ($K\ll 1$). ${\rm Im}(\rho_{10,00}^{\rm ss})$ for three $\tilde{\Omega}$ values at $K=0.1$ (the first column of Fig. 2B), reflecting the functional form of Eq. (20) that maximizes at $\tilde{\Omega}=1$ (or $\Omega=\Omega_{o}$), directly translate to the mean current generated in the 1st qubit ($\langle j_{1}\rangle$, the first column of Fig. 2B). It is also noteworthy that under the same condition ($K=0.1$), the coherence (imaginary part) of the 2nd qubit are effectively zero (${\rm Im}[\rho_{01,00}^{\rm ss}]\approx 0$. See Fig. S1B).
The full expressions for the current, current fluctuations, and density matrix elements of TQS at steady states are too extensive to be shown here explicitly. However, either by noting that the photon current from each qubit is physically generated via two distinct channels of transitions or by using the general expression of the current (see Eq. (61)), $\langle j_{1}\rangle$ and $\langle j_{2}\rangle$ are written in terms of the difference between the steady state populations,
@@
Note that Eq. (28), for instance, expresses the two channels of photon emission from the 1st qubit while the 2nd qubit is in its excited and ground state. In fact, one can also relate the currents with the imaginary parts of coherences, establishing the current-coherence relations for TQS as follows (see Eqs. (S2) – (S5) in Methods for the detailed derivation):
@@
The physical setup of TQS – only the 1st qubit is directly irradiated by the external field, whereas the 2nd qubit is indirectly influenced by the external field through the coupling with the 1st qubit – is reflected in the expressions of Eqs. (30), (31), and (32). The total current from the TQS (Eq. (30)) is similar to the current-coherence relation for the TLS (Eq. (21)) in that it depends only on the imaginary parts of the coherences of the 1st qubit, ${\rm Im}[\rho_{11,01}^{\rm ss}]$ and ${\rm Im}[\rho_{10,00}^{\rm ss}]$, with dominant contributions from ${\rm Im}[\rho_{10,00}^{\rm ss}]$. The current from the 2nd qubit, $\langle j_{2}\rangle$, is determined solely by the correlated coherence, $\rm{Im}\left[\rho^{\rm{ss}}_{10,01}\right]$, with the coupling strength $K$ as the proportionality constant. Notably, the even centrosymmetries of $\langle j_{1,2}\rangle$ with respect to $\delta\omega_{1,2}$ demonstrated in Fig. 2D, i.e., $\langle j_{1,2}\rangle(\delta\omega_{1},\delta\omega_{2})=\langle j_{1,2}\rangle(-\delta\omega_{1},-\delta\omega_{2})$ are consistent only with those of ${\rm Im}[\rho_{11,01}^{\rm ss}]$, ${\rm Im}[\rho_{10,00}^{\rm ss}]$, and ${\rm Im}[\rho_{10,01}^{\rm ss}]$.

### v2

@@ Abstract
The minimal bound of the thermodynamic uncertainty relation (TUR) is modulated from that of the classical counterpart ($\mathcal{Q}_{\rm min}=2$) when a quantumness is present in the dynamical process far from equilibrium. A recent study on a dissipative two-level system (TLS) subject to an external field indicates that quantum coherence can suppress the fluctuations of the irreversible current and loosens the TUR bound to $\mathcal{Q}_{\rm min}^{\rm TLS}\approx 1.25$. Here, we extend on the field-driven single TLS to a quantum-mechanically coupled two-qubit system (TQS), and explore how the quantum coupling between the two qubits, an additional complexity introduced to the probem of TLS, affects the photon current, fluctuations, and the TUR bound. We find that the TUR bound of TQS depends on the strength of coupling, such that $\mathcal{Q}_{\rm min}^{\rm TQS}=\mathcal{Q}_{\rm min}^{\rm TLS}\approx 1.25$ when the two qubits are effectively decoupled under weak coupling, whereas another loose bound $\mathcal{Q}_{\rm min}^{\rm TQS}\approx 1.36$ is identified for two strongly coupled qubits under strong fields. By contrasting the TQS against two coupled noisy oscillators, we illuminate the quantumness unique to the TQS and its effect on the TUR. Our findings from the study of TQS form the basis for understanding the TUR of more general $N$-qubit systems.
@@
Since its first proposal [1] and subsequent proof for continuous time Markov jump processes on networks and overdamped Langevin dynamics [2, 3], the thermodynamic uncertainty relation (TUR) has been extended to more general contexts along with its variations over the past decade [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The TURs are written in the form of an inequality of a product ($\mathcal{Q}$) between total entropy production, $\Delta S_{tot}(t)/k_{B}$, and the squared relative error of a current-like dynamical observable $\theta(t)$ with odd parity $\theta(-t)=-\theta(t)$, satisfying
$\mathcal{Q}=\frac{\Delta S_{tot}(t)}{k_{B}}\frac{\langle\delta\theta^{2}(t)\rangle}{\langle\theta(t)\rangle^{2}}\geq\mathcal{Q}_{\rm min}.$ (1)
This signifies the trade-off between the dissipation and the precision of a dynamical process, with their uncertainty product $\mathcal{Q}$ being lower-bounded. If one were to increase the precision of the process, one should obtain a longer time trace, which requires greater amount of energetic cost. The relation (Eq. (1)) can also be interpreted as a quantitative version of thermodynamic principle for nonequilibrium processes that constrains the total entropy production more tightly than the second law as $\Delta S_{tot}(t)/k_{B}\geq\mathcal{Q}_{\rm min}\langle\theta(t)\rangle^{2}/\langle\delta\theta^{2}(t)\rangle$ [7, 23]. The minimal bound is $\mathcal{Q}_{\rm min}=2$, for Markov jump process and overdamped Langevin systems under constant driving, which encompass a number of conventional situations in nonequilibrium steady states; however, the value of $\mathcal{Q}_{\rm min}$ further reduces for more general nonequilibrium processes, when the system of interest is underdamped [21, 24, 25] or subject to time-dependent drivings [11, 26].
Several studies that examined the TURs for open quantum processes have indicated a loosening of its minimal bound $\mathcal{Q}_{\rm min}<2$, leading to the TUR violation [27, 28, 29, 30, 31, 32, 33, 34, 35], signifying that quantumness can enhance the precision of the dynamical processes beyond its classical limit set by the conventional TUR. Specifically, studies of molecular junctions, made of a single or two quantum dots between the baths of thermal or chemical gradient driving heat or charge transport, examined the TUR or the upper-bound of engine efficiency, induced by the quantum coherence or particle correlation [36, 27, 28, 37, 29, 38, 39, 40].
@@
Our recent works of a field-driven two-level system (TLS) and three-level $\Lambda$-system immersed in photonic bath [41, 42], focusing on the effect of quantumness accompanied with light-matter interaction on the TUR, have shown that the size of the fluctuations in photon current is decided by the competition between the real and imaginary parts of the quantum coherence, corresponding to the off-diagonal elements of the density matrix, that arises from a superposition between different states in a given basis [43]. The imaginary part of the coherence, which was shown to be responsible for the generation of non-equilibrium current in open quantum systems [44, 45, 46], is linked to the dissipative flow of energy or absorption between states that suppresses the current fluctuations. In contrast, the real part, related to the dephasing mechanism accompanied by population mixing, amplifies the fluctuations [46, 41, 42].
@@
Here, we consider a straightforward extension of the one-qubit TLS, which could be related to a variety of physical platforms involving coherent driving of multipartite quantum systems, including not only the above-mentioned molecular junctions [36, 27, 28, 37, 29, 38, 39], but also cold atoms in optical lattices [47], circuit quantum electrodynamics systems [48, 49], and quantum dots in cavity [50, 51]. Specifically, we study a theoretical model of quantum-mechanically coupled two-qubit system (TQS), in which two atoms with ground and excited states [52] are in a photonic bath while an external field continuously irradiates only one of the two atoms. The dynamics in the first atom can affect the quantum state of the second atom due to the coupling, or it simply emits and dissipates a photon into the bath, the process of which can be formulated with the corresponding Lindblad equations.
@@
We characterize non-equilibrium responses of the TQS by means of the uncertainty product of TUR ($\mathcal{Q}$), finding that the maximal precision in the photon current is attained when both the strengths of field and two-qubit coupling are moderate. Our study offers physical insights into this finding by discussing the details of how the coherence, correlation, and entanglement of TQS arise from the varying strength of external field and quantum coupling of the two qubits, and contribute to its TUR.
@@ II.1 Two-qubit system subject to an external field and its evolution equation
where $K$ is the coupling strength. The external field with the frequency $\omega$ irradiates the 1st qubit, resulting in an interaction energy Hamiltonian, is modeled semi-classically as
@@
where the dipole operator $\vec{d}=\vec{d}_{10}\sigma^{1}_{+}+\vec{d}_{01}\sigma^{1}_{-}$ acts on the first qubit and the electric field is given by $\vec{E}(r,t)\simeq\vec{\epsilon}e^{i\omega t}+\vec{\epsilon}^{\ast}e^{-i\omega t}$. In evaluating the effect of $H_{\rm ext}(t)$ on the density matrix, we consider the dipole approximation that the driving field is nearly constant over the scale of our interest, such that $e^{ik\cdot r}\simeq 1$ [53]. Thus, the 1st qubit display unitary oscillation between the ground and the excited states with the Rabi frequency $\Omega=\vec{d}_{10}\cdot\vec{\epsilon}/\hbar=\vec{d}_{01}\cdot\vec{\epsilon}^{\ast}/\hbar$. The Hamiltonian for the surrounding bath is
@@
The Lindblad equation that casts the system-bath interaction in a specific form is obtained by tracing out the bath degrees of freedom and approximating the hierarchical equation under the weak-system bath coupling [54]. The reduced density operator for the system $\rho(t)$ undergoes a dynamic evolution, obeying the following equation along with the Lindblad dissipator, $\mathcal{D}\left(\rho(t)\right)$, acting on both qubits
@@
for $i=1,2$. Here, the term $\bar{n}_{i}=(e^{\beta\hbar\omega_{i}}-1)^{-1}$ denotes the mean occupation number of the bosonic bath affected by the dynamics of the $i$-th qubit, $\{A,B\}=AB+BA$ is the anti-commutator, and $\gamma_{1,2}$ is the relaxation rate of the qubit to its ground state caused by the interaction with the bath. The condition of the weak system-bath coupling is dictated by the condition of $\omega\gg\gamma_{1},\gamma_{2}$. Rescaling the time scale by $\gamma_{1}$ as $t\rightarrow\tau=\gamma_{1}t$, and all the frequencies discussed here as $\gamma_{2}\rightarrow\gamma\equiv\gamma_{2}/\gamma_{1}$, $\Omega\rightarrow\Omega/\gamma_{1}$, and $K\rightarrow K/\gamma_{1}$ yield a set of coupled equations for the density matrix elements.
@@
After applying the rotating-wave approximation (RWA) under the condition of $\omega\gg\Omega$ along with $\omega\gg\delta\omega$, $K$, $\gamma_{1}$ and representing each density matrix element in a rotating frame, we obtain the Liouville equation for the vectorized form of the reduced density matrix in the Fock-Liouville space,
@@ II.3 Uncertainty product for TQS
Using the definition of photon current through the $i$-th qubit $\langle j_{i}\rangle=\langle n_{i}(\tau)\rangle/\tau$, one can recast Eq. (15) into the product between the affinity (entropy production per photon emission, $\mathcal{A}=(\Delta S_{tot}/k_{B})/\langle n(\tau)\rangle$) and the Fano factor ($\mathcal{F}$) of photon current:
@@
where $j=\sum_{i=1,2}j_{i}$. Here, we use the fact that the total entropy production from the system is contributed by the net number of photon emissions that have occurred from both qubits, namely, $\Sigma^{\rm ss}\tau=\mathcal{A}_{1}\langle n_{1}(\tau)\rangle+\mathcal{A}_{2}\langle n_{2}(\tau)\rangle$, with $\mathcal{A}_{i}=\hbar\omega_{i}/k_{B}T$, which thus yields $\mathcal{A}=\frac{\sum_{i=1,2}\mathcal{A}_{i}\langle j_{i}\rangle}{\sum_{i=1,2}\langle j_{i}\rangle}$. The mean current $(\left\langle j\right\rangle)$ and the current fluctuations $({\rm Var}(j))$ at steady states are calculated using the method of cumulant generating function (see Methods).
@@ III.1 Populations and coherences of TQS and current-coherence relations
Fig. 2 demonstrates the coherences and currents of the TQS under varying amount of detunings ($\delta\omega_{1}$, $\delta\omega_{2}$) for select values of Rabi frequency ($\Omega/\Omega_{o}=0.1$, 1, 10) and coupling strength ($K=0.1,1,10$). For the sake of our discussion, we normalize the Rabi frequency by $\Omega_{o}=0.37$ that minimizes $\mathcal{Q}^{\rm TLS}$ (Eq. (23)) [41], defining $\tilde{\Omega}\equiv\Omega/\Omega_{o}$. As commented in Eq. (22), the photon current generated in an isolated qubit interacting with an external field is directly proportional to the imaginary part of coherence [44, 45, 46]. This situation arises when the 1st qubit is subject to the external field, whereas the 2nd qubit is effectively decoupled from the 1st qubit due to the weak coupling ($K\ll 1$). ${\rm Im}(\rho_{10,00}^{\rm ss})$ for three $\tilde{\Omega}$ values at $K=0.1$ (the first column of Fig. 2B), reflecting the functional form of Eq. (21) that maximizes at $\tilde{\Omega}=1$ (or $\Omega=\Omega_{o}$), directly translate to the mean current generated in the 1st qubit ($\langle j_{1}\rangle$, the first column of Fig. 2B). It is also noteworthy that under the same condition ($K=0.1$), the coherence (imaginary part) of the 2nd qubit is effectively zero (${\rm Im}[\rho_{01,00}^{\rm ss}]\approx 0$. See Fig. S1B).
The full expressions for the current, current fluctuations, and density matrix elements of TQS at steady states are too extensive to be shown here explicitly. However, either by noting that the photon current from each qubit is physically generated via two distinct channels of transitions or by using the general expression of the current (see Eq. (66) derived in Methods), $\langle j_{1}\rangle$ and $\langle j_{2}\rangle$ are written in terms of the difference between the steady state populations,
@@
Note that Eq. (29), for instance, expresses the two channels of photon emission from the 1st qubit while the 2nd qubit is in its excited and ground state. In fact, one can also relate the currents with the imaginary parts of coherences, establishing the current-coherence relations for TQS as follows (see Eqs. (S2) – (S5) in SI for the detailed derivation):
@@
The physical setup of TQS – only the 1st qubit is directly irradiated by the external field, whereas the 2nd qubit is indirectly influenced by the external field through the coupling with the 1st qubit – is reflected in the expressions of Eqs. (31), (32), and (33). The total current from the TQS (Eq. (31)) is similar to the current-coherence relation for the TLS (Eq. (22)) in that it depends only on the imaginary parts of the coherences of the 1st qubit, ${\rm Im}[\rho_{11,01}^{\rm ss}]$ and ${\rm Im}[\rho_{10,00}^{\rm ss}]$, with dominant contributions from ${\rm Im}[\rho_{10,00}^{\rm ss}]$. The current from the 2nd qubit, $\langle j_{2}\rangle$, is determined solely by the correlated coherence, $\rm{Im}\left[\rho^{\rm{ss}}_{10,01}\right]$, with the coupling strength $K$ as the proportionality constant. Notably, the even centrosymmetries of $\langle j_{1,2}\rangle$ with respect to $\delta\omega_{1,2}$ demonstrated in Fig. 2D, i.e., $\langle j_{1,2}\rangle(\delta\omega_{1},\delta\omega_{2})=\langle j_{1,2}\rangle(-\delta\omega_{1},-\delta\omega_{2})$ are consistent only with those of ${\rm Im}[\rho_{11,01}^{\rm ss}]$ (Fig. 2A), ${\rm Im}[\rho_{10,00}^{\rm ss}]$ (Fig. 2B), and ${\rm Im}[\rho_{10,01}^{\rm ss}]$ (Fig. 2C).